# Hyp-OW: Exploiting Hierarchical Structure Learning with Hyperbolic Distance Enhances Open World Object Detection

### v1

@@ A.1 Dataset Composition
In this section, we provide the composition of each dataset split and explain the main differences between them. The splits are made with images taken from the PASCAL-VOC2007/2012 (Everingham et al., 2010) and MS-COCO dataset (Lin et al., 2014), containing a total of 80 classes grouped into 12 catogories (SuperClass): vehicle, outdoor, animal, accessory, sports, kitchen, food, furniture, appliance, electronic, indoor, and person. We categorize the different datasets into three regimes based on a semantic similarity metric (defined below), which quantifies the overlap between known and unknown items:
@@
  • High regime : Hierarchical Split
The low regime implies no semantic similarity between classes of different tasks. For instance, OWDETR Split’s ( 3) classes of task 1 are composed of vehicle, animal, and person, which do not appear in tasks 2, 3, or 4. On the other hand, high regime shows high similarity between classes across all tasks. Each task of the Hierarchical split (Table 5) contains at least one class from each of the 12 categories. The OWOD Split is a trade-off between the aforementioned datasets.
@@
Medium regime : OWOD Split
Task 1 Task 2 Task 3 Task 4
aeroplane truck frisbee bed
bicycle traffic light skis toilet
bird fire hydrant snowboard laptop
boat stop sign sports ball mouse
bottle parking meter kite remote
bus bench baseball bat keyboard
car elephant baseball glove cell phone
cat bear skateboard book
chair zebra surfboard clock
cow giraffe tennis racket vase
diningtable backpack banana scissors
dog umbrella apple teddy bear
horse handbag sandwhich hair drier
motorbike tie hot dog toothbrush
person suitcase broccoli wine glass
pottedplant microwave carrot cup
sheep oven hot dog fork
sofa toaster pizza knife
train sink donut spoon
tvmonitor refrigerator cake bowl
High regime : hierarchical Split
Task 1 Task 2 Task 3 Task 4
bicycle person bus truck
car motorbike train boat
traffic light aeroplane parking meter bench
fire hydrant stop sign cow zebra
bird horse elephant giraffe
cat sheep bear tie
dog umbrella handbag suitcase
backpack snowboard kite skateboard
frisbee sports ball baseball bat surfboard
skis cup baseball glove tennis racket
bottle sandwich fork spoon
wine glass orange knife bowl
banana broccoli carrot pizza
apple pottedplant hot dog donut
chair bed diningtable cake
sofa laptop remote toilet
tvmonitor mouse keyboard cell phone
microwave toaster sink refrigerator
oven clock scissors hair drier
book vase teddy bear toothbrush
@@
We propose a measure to quantify the semantic similarity overlap between known and unknown classes in each dataset. To achieve this, we utilize the GloVe (Pennington et al., 2014) embedding for each known class $c\in\mathcal{K}^{t}$ (and unknown class $k\in\mathcal{U}^{t}$), denoted by $\mathbf{\omega_{c}}$ ( respectively $\mathbf{\omega_{k}}$ ). For a given task $t\in[1,T]$, we define the semantic overlap between knowns and unknowns as:
$S_{t}=\frac{1}{|\mathcal{U}^{t}|}\displaystyle{\sum_{k\in\mathcal{U}^{t}}\max_{c\in\mathcal{K}^{t}}}\frac{<\mathbf{\omega^{c}},\mathbf{\omega^{k}>}}{\lVert\mathbf{\omega^{c}}\rVert_{2}\cdot\lVert\mathbf{\omega^{k}}\rVert_{2}},\forall t\leq T-1$ (8)
Higher value means higher similarity between the knowns and unknowns. For a given task $t$, this value calculates the similarity between each unknown and the highest-ranked known and averages it across all unknowns. Figure 5 shows the evolution of this similarity measure throughout the training. As expected, the three splits ranks monotonically as low similarity (OWDETR Split ), medium similarity (OWOD Split) and high similartity (Our hierarchical Split).
@@
Appendix B Implementation Details
@@
Appendix C Additional Experimental Results
C.1 Detailed Results
@@
Hyp-OW (Ours) 23.9 (+6.3) 72.7 23.3 (+1.0) 59.8 42.2 50.6 25.4 49.3 39.8 46.2 (+4.2) 46.4 40.3 44.8 (+4.9)
@@
Hyp-OW (Ours) 23.5 $\boldsymbol{(+4.1)}$ 59.4 20.6 $\boldsymbol{(+3.1)}$ 51.4 37.4 44.4 26.3 $\boldsymbol{(+6.6)}$ 42.9 24.6 36.8 37.4 22.4 33.6 (+2.1)
@@
Hyp-OW (Ours) 34.9 (+5.2) 49.9 47.5 $\boldsymbol{(+3.5)}$ 42.0 49.0 45.5 $\boldsymbol{(+2.7)}$ 55.2 $\boldsymbol{(+2.5)}$ 44.4 44.1 44.3 $\boldsymbol{(+3.1)}$ 42.8 47.0 43.9 $\boldsymbol{(+3.0)}$
@@
Next, we turn our attention to the issue of unknown object confusion and address it through the Absolute Open-Set Error (A-OSE) metric, as introduced by Joseph et al. (2021). The A-OSE quantifies the number of unknown objects that are incorrectly classified as known (lower is better). In Table 9, we present the A-OSE values alongside the U-Recall metric for the Hierarchical Split.
hierarchical Split Task 1 Task 2 Task 3
U-Recall ($\uparrow$) A-OSE($\downarrow$) U-Recall ($\uparrow$) A-OSE($\downarrow$) U-Recall ($\uparrow$) A-OSE($\downarrow$)
@@
Table 9: Unknown Confusion. Hyp-OW achieves the highest U-Recall, indicating its superior ability to detect unknown objects. Additionally, it exhibits the lowest A-OSE, further emphasizing its capability to effectively differentiate between known and unknown objects.
C.2 Unveiling Co-occurrence Learning: The Influence of Bounding Box Overlap
In the Object Detection (OD) task, where bounding boxes can overlap, the model has a tendency to learn co-occurrence patterns of frequently appearing classes. This phenomenon is quantitatively demonstrated in the Heatmap Figure 6, where the model exhibits high similarity between classes that do not have any semantic relationship, such as ’person’ and ’tie’, or ’teddy bear’ and ’bicycle’ (highlighted in teal). This observation is further supported by Figures  7,  8 and  9, which provide qualitative examples of this co-occurrence learning.
@@ C.2 Unveiling Co-occurrence Learning: The Influence of Bounding Box Overlap
Appendix D Ablation Analysis
@@
D.1 Detailed Ablation Results
@@
w/ Cosine Distance ($c=0$) 32.8 $\boldsymbol{(-2.1)}$ 49.0 $\boldsymbol{(-0.9)}$ 46.4 $\boldsymbol{(-1.1)}$ 42.0 48.8 45.4 55.4 43.6 42.3 43.2(1.$\boldsymbol{(-1.1)}$ 42.4 45.2 43.1
w/o SuperClass Regularizer ($\beta=0$) 32.0 $\boldsymbol{(-2.9)}$ 50.0 47.1 41.3 48.9 45.1 52.9 $\boldsymbol{(-2.3)}$ 43.8 43.5 43.7 42.8 45.9 43.5
w/o Adaptive Relabeling 34.7 41.2 $\boldsymbol{(-8.7)}$ 47.6 36.2 41.5 38.9 $\boldsymbol{(-6.6)}$ 54.1 $\boldsymbol{(-1.1)}$ 36.5 36.6 36.5 $\boldsymbol{(-7.8)}$ 34.5 37.9 36.1 $\boldsymbol{(-7.8)}$
Table 10: Impact of each component of Hyp-OW on Hierachichal Split.
@@
To asses the impact of this module, we adopt PROB’s (Zohar et al., 2022) methodology where every unmatched $\mathbf{q^{\textit{u}}}$ is labelled as unknowns.
@@
D.2 Effectiveness of our Semantic Adaptive Relabeling Scheme
We now provide insights into the effectiveness of our Semantic Distance-based Relabeling Scheme. Figure 10 and  11 show the heatmap distance between known (left) and unknown items (bottom) for Task 1 in the OWOD and Hierarchical Split, respectively. Lighter colors indicate lower distances and higher similarity between known and unknown classes.
@@ D.3 SuperClass Regularizer
D.4 Cosine versus Hyperbolic Distance
@@
D.5 Impact of Curvature Coefficient c
To investigate the influence of the curvature coefficient on the hyperbolicity of the embedding space, we conducted experiments using different values of $c$. Our focus was on Task 1 of the Hierarchical Split, where we assessed the model’s performance for four values of $c$: $0.0$, $0.1$, $0.2$, and $0.5$. The results are presented in Table 11. Notably, the curvature coefficient primarily impacts the U-Recall metric, while the mAP metric remains relatively stable. This observation is expected since the curvature coefficient $c$ directly influences the learned hierarchical structure (hence the Semantic Similarity Distance), which is the cornerstone of our Adaptive Relabeling Scheme for unknown retrieval. Among the tested values, we found that $c=0.1$ yielded the optimal performance, consistent with previous findings in the literature (Ermolov et al., 2022; Khrulkov et al., 2020b; Yue et al., 2023).
Hierarchical Split - Task 1
@@ D.5 Impact of Curvature Coefficient c
Table 11: Impact of curvature coefficient $c$.

### v2

@@ Appendix B Dataset Information
In this section, we provide the composition of each dataset split and explain the main differences between them. The splits are made with images taken from the PASCAL-VOC2007/2012 (Everingham et al. 2010) and MS-COCO dataset (Lin et al. 2014), containing a total of 80 classes grouped into 12 catogories (SuperClass): vehicle, outdoor, animal, accessory, sports, kitchen, food, furniture, appliance, electronic, indoor, and person. We categorize the different datasets into three regimes based on a semantic similarity metric (defined below), which quantifies the overlap between known and unknown items:
  • High regime : Hierarchical Split
  • Medium regime : OWOD Split
@@
The low regime implies no semantic similarity between classes of different tasks. For instance, OWDETR Split’s (Table 8) classes of task 1 are composed of vehicle, animal, and person, which do not appear in tasks 2, 3, or 4. On the other hand, high regime shows high similarity between classes across all tasks. Each task of the Hierarchical split (Table 5) contains at least one class from each of the 12 categories (Table 6 shows the statistics of the split). The OWOD Split is a trade-off between the aforementioned datasets.
High regime : hierarchical Split
Task 1 Task 2 Task 3 Task 4
bicycle person bus truck
car motorbike train boat
traffic light aeroplane parking meter bench
fire hydrant stop sign cow zebra
bird horse elephant giraffe
cat sheep bear tie
dog umbrella handbag suitcase
backpack snowboard kite skateboard
frisbee sports ball baseball bat surfboard
skis cup baseball glove tennis racket
bottle sandwich fork spoon
wine glass orange knife bowl
banana broccoli carrot pizza
apple pottedplant hot dog donut
chair bed diningtable cake
sofa laptop remote toilet
tvmonitor mouse keyboard cell phone
microwave toaster sink refrigerator
oven clock scissors hair drier
book vase teddy bear toothbrush
Task 1 Task 2 Task 3 Task 4
##\## training images 28,677 38,552 27,147 34,136
##\## test images 4,952 4,952 4,952 4,952
##\## unknowns 26,523 10,476 5,441 N/A
Table 6: Statistics of Hierarchical Split.
@@
Task 1 Task 2 Task 3 Task 4
aeroplane truck frisbee bed
bicycle traffic light skis toilet
bird fire hydrant snowboard laptop
boat stop sign sports ball mouse
bottle parking meter kite remote
bus bench baseball bat keyboard
car elephant baseball glove cell phone
cat bear skateboard book
chair zebra surfboard clock
cow giraffe tennis racket vase
diningtable backpack banana scissors
dog umbrella apple teddy bear
horse handbag sandwhich hair drier
motorbike tie hot dog toothbrush
person suitcase broccoli wine glass
pottedplant microwave carrot cup
sheep oven hot dog fork
sofa toaster pizza knife
train sink donut spoon
tvmonitor refrigerator cake bowl
@@
We propose a measure to quantify the semantic similarity overlap between known and unknown classes in each dataset. To achieve this, we utilize the GloVe (Pennington, Socher, and Manning 2014) embedding for each known class $c\in\mathcal{K}^{t}$ (and unknown class $k\in\mathcal{U}^{t}$), denoted by $\mathbf{\omega_{c}}$ ( respectively $\mathbf{\omega_{k}}$ ). For a given task $t\in[1,T-1]$, we define the semantic overlap between knowns and unknowns as:
$S_{t}=\frac{1}{|\mathcal{U}^{t}|}\displaystyle{\sum_{k\in\mathcal{U}^{t}}\max_{c\in\mathcal{K}^{t}}}\frac{<\mathbf{\omega^{c}},\mathbf{\omega^{k}>}}{\lVert\mathbf{\omega^{c}}\rVert\cdot\lVert\mathbf{\omega^{k}}\rVert},\forall t\leq T-1$ (9)
Higher value indicates larger similarity between the knowns and unknowns. Figure 6 shows the evolution of this similarity measure throughout the training. This measure quantifies the similarity overlap between known and unknown items, with higher values indicating larger overlap. The three splits, OW-DETR Split (Low regime), OWOD Split (Medium regime), and Hierarchical Split (High regime), consistently align with our intended design, providing a foundation for evaluating baseline methods across diverse scenarios. Note that this metric is increasing as the number of knowns grows throughout the training.
@@
Appendix C Implementation Details
@@
Parameters Range Comments
$\alpha$$\{0.05,0.1,0.2\}$
$\beta$$\{0.05,0.1,0.2\}$ should be lower than $\alpha$
$\tau_{2}$$\{0.02,0.05,0.1\}$ should be lower than $\tau_{1}$ as in (Li et al. 2023)
$m$$\{10,20,50\}$ Negligible impact, yet increases memory
$c$$\{0.0,0.1,0.2,0.5\}$ See Table 2
$k$$\{1,2,5\}$ Negligible impact, yet increases training time
Table 11: Hyperparameters Tuning Ranges.
@@
Appendix D Additional Experimental Results
Detailed Results
@@ Appendix D Additional Experimental Results
Methods Current Previous Current Both Previous Current Both Previous Current Both
@@
Hyp-OW (Ours) 23.9 72.7 23.3 59.8 42.2 50.6 25.4 49.3 39.8 46.2 46.4 40.3 44.8
$\Delta$(Rel. Difference) +6.3 $\leq 1.0$ +1.0 $\leq 1.0$ $\leq 1.0$ $\boldsymbol{+4.2}$ $\boldsymbol{+4.9}$
@@
Hyp-OW (Ours) 23.5 59.4 20.6 51.4 37.4 44.4 26.3 42.9 24.6 36.8 37.4 22.4 33.6
$\Delta$(Rel. Difference) $\boldsymbol{+4.1}$ $\leq 1.0$ $\boldsymbol{+3.1}$ $\leq 1.0$ $\boldsymbol{+6.6}$ $\leq 1.0$ +2.1
@@
$\Delta$(Rel. Difference) +5.5 $\leq 1.0$ $\boldsymbol{+3.6}$ $\boldsymbol{+2.6}$ $\boldsymbol{+2.5}$ $\boldsymbol{+3.0}$ $\boldsymbol{+2.9}$
@@
Next, we turn our attention to the issue of unknown object confusion and address it through the Absolute Open-Set Error (A-OSE) metric, as introduced by (Joseph et al. 2021). The A-OSE quantifies the number of unknown objects that are incorrectly classified as known (lower is better). In Table 13, we present the A-OSE values alongside the U-Recall metric for the Hierarchical Split.
Regime Task 1 Task 2 Task 3
Methods U-Recall ($\uparrow$) A-OSE($\downarrow$) U-Recall ($\uparrow$) A-OSE($\downarrow$) U-Recall ($\uparrow$) A-OSE($\downarrow$)
Medium OW-DETR 7.5 10,240 6.2 8,441 5.7 6,803
PROB 19.4 5,195 17.4 6,452 19.6 2,641
Hyp-OW (Ours) 23.5 11,275 20.6 4,805 26.3 3,548
High
@@
Table 13: Unknowns Confusion. Hyp-OW achieves the highest U-Recall, indicating its superior ability to detect unknown objects. The misclassification rate (A-OSE) shows improved performance for Hyp-OW in the high regime, where inherent hierarchical structure is more pronounced than the medium regime.
Qualitative Results
In this section, we present a qualitative visualization that clearly illustrates the superior detection precision of our method, Hyp-OW, in comparison to the PROB baselines (Figure 7). Additionally, we provide a detailed quantitative assessment underscoring the impact of each individual component within the Hyp-OW framework (Figure 8).
Figure 7: Unknown detection illustrated for Hyp-OW (first row) and PROB (second row), highlighted with the red boxes. By learning the hierarchical structure of items, Hyp-OW provides a more fine-grained detection of objects. On the other hand, PROB, which learns a generalized objectness score by averaging all features, is restricted to a more broad, coarse-grained detection approach
Figure 8: Impact of each component of Hyp-OW on known items accuracy. Hyp-OW (left column) exhibits not only higher confidence but also superior capability in identifying correct items.
Unveiling Co-occurrence Learning: The Influence of Bounding Box Overlap
In the Object Detection (OD) task, where bounding boxes can overlap, the model has a tendency to learn co-occurrence patterns of frequently appearing classes. This phenomenon is quantitatively demonstrated in the Heatmap Figure 9, where the model exhibits high similarity between classes that do not have any semantic relationship, such as ’person’ and ’tie’, or ’teddy bear’ and ’bicycle’ (highlighted in teal). This observation is further supported by Figures  10,  11 and  12, which provide qualitative examples of this co-occurrence learning.
@@
Appendix E Ablation Analysis
@@
Detailed Ablation Results
@@
Medium
Hyp-OW (Ours) 23.5 59.4 20.6 51.4 37.4 44.4 26.3 42.9 24.6 36.8 37.4 22.4 33.6
with Cosine Distance ($c=0$) 21.7 59.1 19.2 51.6 33.6 42.6 25.5 40.5 23.9 35.0 36.7 20.7 32.7
w/o SuperClass Regularizer ($\beta=0$) 22.6 59.0 19.0 50.4 36.8 43.6 23.7 41.4 24.7 35.8 37.8 22.0 33.8
w/o Adaptive Relabeling 22.7 44.6 18.7 38.5 28.1 33.3 27.2 32.7 40.7 28.8 29.8 19.3 32.1
High
@@
w/ Cosine Distance ($c=0$) 32.8 49.0 46.4 42.0 48.8 45.4 55.4 43.6 42.3 43.2 42.4 45.2 43.1
w/o SuperClass Regularizer ($\beta=0$) 32.0 50.0 47.1 41.3 48.9 45.1 52.9 43.8 43.5 43.7 42.8 45.9 43.5
w/o Adaptive Relabeling 34.7 41.2 47.6 36.2 41.5 38.9 54.1 36.5 36.6 36.5 34.5 37.9 36.1
Table 14: Impact of each component of Hyp-OW on OWOD and Hierachichal Split.
@@
To asses the impact of this module, we adopt PROB’s (Zohar, Wang, and Yeung 2023) methodology where every unmatched $\mathbf{q^{\textit{u}}}$ is labelled as unknowns.
@@
We here demonstrate how Euclidian distance and cosine similarity are linked in case of normalized vectors. Recall that the Euclidian distance is recovered when $c\to 0$: $\lim_{c\to 0}d_{hyp}(\mathbf{x},\mathbf{y})=2\rVert\mathbf{x}-\mathbf{y}\rVert$
For two normalized vectors $\mathbf{x},\mathbf{y}$:
$2^{2}\lVert\frac{\mathbf{x}}{\mathbf{\lVert x\lVert}}-\frac{\mathbf{y}}{\mathbf{\lVert y\lVert}}\lVert^{2}=4(\lVert\frac{\mathbf{x}}{\mathbf{\lVert x\lVert}}\lVert^{2}+\lVert\frac{\mathbf{y}}{\mathbf{\lVert y\lVert}}\lVert^{2}-2\frac{<\mathbf{x},\mathbf{y}>}{\lVert\mathbf{x}\rVert\cdot\lVert\mathbf{y}\rVert})=4(2-2\frac{<\mathbf{x},\mathbf{y}>}{\lVert\mathbf{x}\rVert\cdot\lVert\mathbf{y}\rVert})=4\times d_{cos}(\mathbf{x},\mathbf{y})$ (10)
Effectiveness of our Semantic Adaptive Relabeling Scheme
We now provide insights into the effectiveness of our Semantic Distance-based Relabeling Scheme. Figure 13 and  14 show the heatmap distance between known (left) and unknown items (bottom) for Task 1 in the OWOD and Hierarchical Split, respectively. Lighter colors indicate lower distances and higher similarity between known and unknown classes.
@@
Cosine versus Hyperbolic Distance
@@
Impact of Curvature Coefficient c
To investigate the influence of the curvature coefficient on the hyperbolicity of the embedding space, we conducted experiments using different values of $c$. Our focus was on Task 1 of the Hierarchical Split, where we assessed the model’s performance for four values of $c$: $0.0$, $0.1$, $0.2$, and $0.5$. The results are presented in Table 2. Notably, the curvature coefficient primarily impacts the U-Recall metric, while the mAP metric remains relatively stable. This observation is expected since the curvature coefficient $c$ directly influences the learned hierarchical structure (hence the Semantic Similarity Distance), which is the cornerstone of our Adaptive Relabeling Scheme for unknown retrieval. Among the tested values, we found that $c=0.1$ yielded the optimal performance, consistent with previous findings in the literature (Ermolov et al. 2022; Khrulkov et al. 2020b; Yue et al. 2023).
@@
Table 15: Impact of curvature coefficient $c$ for Hierarchical Split Task 1.